# Information Discovery in e-Commerce

### v1

@@ 7.1.2 Characteristics of e-commerce question answering
E-commerce QA services focus on answering product-aware questions asked by e-commerce users. Early studies on e-commerce QA focused on providing answers automatically from reviews by heuristic methods [li2009answering, moghaddam2011aqa, yu2012answering]. With the development of both QA techniques and e-commerce services, e-commerce QA has received increasing attention in recent years [mcauley2016addressing, yu2017modelling, yu2018aware, fan2019reading, zhang2020answer, gao2019product, gaoshen2021tois, feng2021multi, deng2022toward].
In contrast with open-domain QA, distinct characteristics of e-commerce QA exist: (i) Domain-specific aspects are the first e-commerce QA characteristic. E-commerce QA systems rely on exploiting more domain-specific information from product descriptions. Different products make product-aware aspects popular in e-commerce QA studies [mcauley2016addressing]. These product-aware aspects can help distinguish products and questions. (ii) The second characteristic of e-commerce QA is the large number of consumer reviews, which can be used as a data source to help people form opinions and decisions [liu2016retrieving, mcauley2016addressing]. With the growth of those opinionated reviews, e-commerce users rely on advice from reviews before making purchase decisions. Reviews have been used as supporting data and candidate answers to supervise QA prediction models [yu2018aware]. (iii) The third characteristic of e-commerce QA is various sources of answers. Most e-commerce QA services focus on extracting answers from reviews [mcauley2016addressing], and many e-commerce sites provide question answer pairs as knowledge bases for QA.
Text generation approaches have been studied to generate answers to given questions and reviews. Besides given question answers in many e-commerce QA platforms, question reranking and answer reranking also have been studied [yu2017modelling, zhang2020answer]. Existing e-commerce QA research can be divided into two directions: extractive product-aware QA and generative product-aware QA. The former focuses on extracting sentences or passages from reviews to answer questions, whereas the latter applies textual generation approaches to generate answers. We detail each type of e-commerce QA study in Section 7.1.3 and Section 7.1.4 respectively.
@@ 7.1.3 Extractive product-aware QA
Most existing e-commerce QA systems aim at extracting relevant sentences or fragments from the input text to answer the question given by the consumer. Early studies automatically extracted answers from reviews by heuristic unsupervised methods [li2009answering, moghaddam2011aqa], whereas the follow-ups mainly focused on the matching between questions and reviews or candidate answers [mcauley2016addressing, yu2018aware]. yu2012answering proposed a framework for opinionated QA, which organizes reviews into a hierarchy structure and retrieves review sentences as the answer. The authors then used such a hierarchy structure to help retrieve question analysis and relevant review fragments. A joint optimization approach was proposed by simultaneously considering review salience, coherence, and diversity to rank fragments. liu2016retrieving aimed to find a concise set of questions addressed by a given review and cover its main points to help the user quickly comprehend the reviews. The authors proposed a two-stage framework, where a probabilistic retrieval model was used to retrieve candidate questions first. Subsequently, a matching procedure between answers and questions was used to bridge the vocabulary gap between reviews and questions. Then, increasing extractive QA strategies were proposed to tackle challenges in e-commerce QA.
@@
Unlike open-domain QA in other scenarios that require utilizing entity knowledge graphs, any two products in e-commerce sites are distinct even though they belong to the same category. Some products such as clothes and paintings do not even have proper names. Hence aspects have been considered to replace the external knowledge of e-commerce to address this problem. mcauley2016addressing proposed an answer prediction model by incorporating an aspect analytic model to learn latent aspect-specific review representation for predicting the answer. wan2016modeling aimed to address ambiguity, subjectivity, and diversity problems in consumer reviews. By leveraging multiple answers in a supervised framework, the authors aim to provide more accurate answers to those objective and subjective questions. The authors also released a large-scale e-commerce QA dataset consisting of 135 thousand products from Amazon, 808 thousand questions, 3 million answers, and 11 million reviews. carmel2018product focused on subjective questions from Amazon customers, which can relate to various intent types such as product usage, recommendations, and opinions. The authors applied automatic QA methods, enhanced with community QA approaches to retrieve the most relevant answer found in reviews and QAs to address this problem. yu2018aware proposed an answer prediction model by incorporating an aspect analytic model to learn latent aspect-specific review representation for predicting the answer. The authors have verified its advantage of generating aspect-specific representations for new questions, which they used to develop the predictive answer model to capture intricate relationships among question texts and review texts jointly. The proposed model uses reviews as knowledge to predict the answer by classifying answers into two types, binary (i.e. “yes” or “no”) and open-ended responses. As the amount of labeling data is limited in customer reviews, das2019learning proposed an adversarial review-based approach to answer subjective and specific product-aware questions in a weakly supervised setting. Reading comprehension has recently been verified as useful to help extract relevant answers from e-commerce reviews [fan2019reading, xu2019review, zhang2020answering, chen2019answer]. Fed by the raw text of product-aware questions and customer reviews, fan2019reading proposed an end-to-end neural network model to synthesize multiple review representations. chen2019answer proposed a multi-task attentive model, namely QAR-Net, to identify plausible answers from product reviews for user questions. QAR-Net can leverage generated question answer pairs to help question-review matching.
@@
Pre-trained language models help understand the content of questions and reviews. xu2019review applied a BERT-based fine-tuning approach to extract answers from reviews. mittal2021distantly utilized a pre-trained language model to learn a vital relevance function by jointly learning unified syntactic and semantic representations of questions and reviews. A QA dataset for review comprehension with subjectivity labels for questions and answers has also been exploited [bjerva2020subjqa]. Besides user reviews, another kind of information, namely product details provided by the manufacturer, has been considered an auxiliary information source for addressing product-related questions [zhang2020answering]. In order to alleviate the unavailability of labeled data, DBLP:conf/emnlp/JainRA23 proposed a distant supervision based NLI model to prepare training data without any manual effort. The end-to-end manner and an auxiliary contrastive loss for evidence extraction are introduced in the training stage.
Review-based QA approaches extract answers from customer reviews, which can partially address these questions. However, there are many products with few or no reviews available. By collecting question answer pairs from real users, many e-commerce platforms develop retrieval-based QA systems for automatically answering frequently asked questions (FAQs) in the e-commerce industry [yu2017modelling, song2020tcnn, song2021online, zhang2020answer]. As shown in Fig. 7.1, we see the retrieval-based QA framework applied in Alibaba, the largest e-commerce platform in China [yu2017modelling]. Given a large scale of question answer pairs (i.e., the knowledge base in Fig. 7.1), a key component is the question rerank module, which reranks candidate questions in a question answering knowledge base to find the best match given a question from a user. Based on such a framework, yu2017modelling formulated e-commerce QA as a paraphrase identification problem, where the target is to identify semantic relations of the given sentence pairs. The authors proposed a transfer learning QA strategy to adapt the shared knowledge learned from a resource-rich source domain to a resource-poor target domain. Amazon presented a large review-based QA dataset, namely AmazonQA, based on their real-world community QA platform [gupta2019amazonqa]. AmaonQA uses consumer reviews as the data resource and extracts snippets to answer questions. song2020tcnn improved the matching performance in retrieval-based e-commerce QA by introducing a multi-layer triple convolutional neural network model. Also, a sub-graph searching mechanism is verified to improve the efficiency of retrieval-based e-commerce QA [song2021online]. zhang2020answer focused on answer selection in retrieval-based e-commerce QA. Using graph neural networks, the authors proposed a new framework to jointly model multiple semantic relations, including the semantic relevance between the question and answers, the textual similarity among answers, and the textual entailment between answers and reviews. rozen2021answering proposed an answer prediction approach that leverages similar questions about other products. The authors calculated the contextual product similarity to determine whether two products are similar in the context of a specific question. Two large-scale datasets, including a question-to-question similarity dataset from Amazon and a corpus of question answer pairs from Amazon, are released with the publication.
@@ 7.1.4 Generative product-aware QA
Many e-commerce portals have provided question answering services that assist users in posing product-aware questions to other consumers who have purchased the same product before. Under the e-commerce QA circumstances, users must read the product’s reviews to find the answer themselves. Given product attributes and reviews, following a cascading procedure, an answer is manually generated. (i) A user skims reviews and finds relevant sentences; (ii) they extract functional semantic units; and (iii) and the user jointly combines these semantic units with attributes and writes an appropriate answer. However, the information overload phenomenon makes this procedure an energy-draining process to pursue an answer from a rapidly increasing number of reviews [gao2019product]. Several strategies have been proposed to automatically generate answers using the product’s reviews to alleviate the burdens of customers [gao2019product, chen2019driven, deng2020opinion, lu2020chime, feng2021multi, deng2022toward]. The task on which these approaches focus is the generative product-aware QA given reviews and product attributes.
@@
In their first attempt, gao2019product, chen2019driven proposed the task of product-aware answer generation, where a product-related question answering model is applied to incorporate customer reviews with product attributes. The authors formulated the research problem in generative e-commerce QA: for a product, there is a question $X^{q}=\{x^{q}_{1},x^{q}_{2},\dots,x^{q}_{T_{q}}\}$, $T_{r}$ reviews $X^{r}=\{x^{r}_{1},x^{r}_{2},\dots,x^{r}_{T_{r}}\}$ and $T_{a}$ key-value pairs of attributes $A=\{(a^{k}_{1},a^{v}_{1}),(a^{k}_{2},a^{v}_{2}),\dots,(a^{k}_{T_{a}},a^{v}_{T_{a}})\}$, where $a^{k}_{i}$ is the name of $i$-th attribute and $a^{v}_{i}$ is the attribute content. Each attribute, including key $a^{k}_{i}$ and value $a^{v}_{i}$, is represented as a single word in the generation task. Given a question $X^{q}$, an answer generator reads the reviews $X^{r}$ and attributes $A$, then generates an answer $\hat{Y}=\{\hat{y}_{1},\hat{y}_{2},\dots,\hat{y}_{T_{y}}\}$. The goal is to generate an answer $\hat{Y}$ that is grammatically correct and consistent with product attributes and opinions in the reviews.
@@
As shown in Fig. 7.2, gao2019product provided an overview of their proposed product-aware answer generator, the PAAG model. PAAG is divided into four parts: (i) review readerreads the review to extract relevant semantic parts; (ii) attribute encoderencodes the attribute key-value pairs using a key-value memory network; (iii) facts decodergenerates the final answer according to the facts learned by the two modules introduced before; and (iv) consistency discriminatordistinguishes whether the generated answer matches the extracted facts, and we also use the result of the discriminator as another training signal. A generative e-commerce QA dataset extracted from JD.com is released with the publication. Similarly, chen2019driven proposed a noise-tolerant solution based on convolutional neural networks to generate natural answers.
@@
It is necessary to consider the text information from different reviews and attributes to answer specific questions in the wild. In Fig. 7.3feng2021multi provided examples to demonstrate the multi-type text relation for product-aware question answering. As an example, Q1 asks “Does the design of this top look baggy?”. R1 and R2 do not answer this question directly. But they provide a common entity “bat-like sleeve”. If we transfer the information provided by R1 and R2 to answer Q1 indirectly, it is easy to generate the answer that “The design of this tops looks baggy.”. The above example shows that we may generate more accurate and pleasing answers to complex questions by integrating, understanding, and reasoning over the information of reviews and product attributes.
@@
One major limitation of most existing generative QA approaches is that they analyze each review and the corresponding attribute of the product individually, i.e., they neglect the relationship between different reviews/attributes of the product. feng2021multi proposed a review-attribute heterogeneous graph neural network, namely RAHGNN, for product-aware answer generation to sufficiently understand and reason the related information and its inner logic in multi-type texts. Most generative product-aware QA methods neglect personalization as it is insufficient to provide the same “completely summarized” answer to all customers. deng2022toward proposed a personalized answer generation method, namely PAGE, to comprehensively model multi-perspective user preferences in personalized product question answering.
@@ 7.2 Dialogue systems in e-commerce
Dialogue systems have attracted more and more attention in e-commerce. This section introduces studies on dialogue systems that can be applied to e-commerce platforms. According to most previous work investigating this research problem Chen2017A, ziwwsdm1, we divide this section into three parts: we briefly introduce recent studies on dialogue systems in Section 7.2.1; then detail task-oriented dialogue systems in e-commerce in Section 7.2.2; demonstrate knowledge-grounded conversational agents in Section 7.2.3.
@@ 7.2.1 Introduction to dialogue systems
In recent years, dialogue systems have received more and more attention in numerous applications, from e-commerce technical support to personal assistant tools [song2017summarizing, Chen2017A, ziwwsdm1, sun2016conversational, zhang2018towards, lei2018sequicity, liu2018knowledge, meng2020dukenet, sun2021conversations, shen2021vida, zhao2021jddc, liu2021conversational, ren2022variational, li2022knowledge, yu2022xdai]. The goal of creating an automatic human-computer conversation system as our assistant or chat companion is no longer an illusion as two different aspects have been combined simultaneously. First, many conversation logs are accessible, making it possible to learn how to respond to the input utterances. Second, deep generative neural network models, such as sequence-to-sequence and generative adversarial networks, have been verified to capture complex patterns in big data [Chen2017A]. Based on these two aspects, studies on dialogue systems aim to provide a natural and coherent response given an utterance from a user [Young2013POMDP, Ritter2011Data, Banchs2013IRIS, Ameixa2014Luke].
Dialogue systems can be divided into chitchat systems, task-oriented dialogue systems, and knowledge-grounded conversations [Chen2017A]. Chitchat agents are applied widely in open-domain dialogue systems, where dialogue systems interact with humans to provide reasonable and natural responses for open-domain dialogues [ziwwsdm1, yan2017building]. Chitchat messages usually represent user experiences and preferences, playing an essential role in many real-world applications. yan2017building revealed most utterances in the online shopping scenario are chitchat messages.
Task-oriented dialogue systems aim to complete a specific task, e.g., restaurant reservation, along with a response generation process. Fig. 7.4 shows traditional task-oriented dialogue systems based on four individual modules: natural language understanding, dialogue state tracking, policy learning, and natural language generation [wen2017latent, mrkvsic2015multi]. Given an utterance from a user, the system will generate a proper response to address the user’s intention. In recent years, end-to-end task-oriented dialogue generation methods have been proposed to address the research problem more efficiently [wen2017network, lei2018sequicity, wu2019transferable].
@@
Knowledge-grounded conversations focus on generating a response with the correct knowledge to address the user’s utterance [meng2020dukenet]. According to the knowledge modality, existing works for knowledge-grounded conversation can be mainly categorized into two groups. The methods in the first group aim to leverage structured knowledge (given knowledge graphs)  [zhou2018commonsense, liu2018knowledge, tuan2019dykgchat, wu2019proactive, moon2019opendialkg, wu2020diverse, zhou-etal-2020-kdconv, wang2020improving, wu2020improving, xu2020knowledge, xu2020conversational, jung2020attnio, xu-etal-2021-discovering]. Those in the second group focus on leveraging unstructured knowledge, such as document-based unstructured knowledge (given a whole document, e.g., a Wikipedia article) [chuanmeng2020refnet, ma2020survey, ma2020compare, tian-etal-2020-response, ren2019thinking, gopalakrishnan2019topical, zekangli2018incremental, qin-etal-2019-conversing, moghe2018towards, zhou2018dataset] or piece-based unstructured knowledge (given some separate pieces of knowledge, e.g., Foursquare tips) [ghazvininejad2018knowledge, dinan2018wizard, meng2020dukenet, kim2020sequential, Lian2019Learning, zheng2019enhancing, zheng2020difference, chen2020bridging, zheng2020approximation, zhao2020knowledge, yu2022xdai].
@@
According to the way of generating responses, dialogue systems can be divided into retrieval-based and generation-based dialogue systems. The former retrieves several response candidates from a prebuilt index and then selects an appropriate one as a response. In contrast, the latter directly synthesizes a reply via natural language generation techniques [hvred, tao2021survey].
@@
Retrieval-based dialogue systems retrieve several response candidates from a prebuilt index and then select an appropriate one as a response. The flourishing of social networking services has accumulated a significant amount of conversation data among humans on the web, encouraging researchers to investigate data-driven approaches to re-use the existing human conversations and select a response for new input from candidates [tao2021survey, xu2021response]. Retrieval-based dialogue generation methods are superior to their generation-based counterparts in response fluency and informativeness, which powers a series of real-world applications, e.g., XiaoIce from Microsoft [zhou2020design]. Learning to rank and matching approaches have been widely applied in the retrieval process [yan2016learning, wu2017sequential, zhou2018multi, yang2018response, yuan2019multi, su2020dialogue, tao2021survey, lin2022task]. A core task in retrieval-based dialogue systems is response selection. According to the way of matching candidate responses, existing studies on retrieval-based response selection can be divided into $3$ types: representation-based, interaction-based, and pre-trained language model-based methods. Representation-based methods are composed of a representation-matching paradigm and consist of a representation layer and a matching layer [yan2016learning, wu2018response, wang2017deep, zhou2018dataset, zhou2018multi, yan2018response, xu2021topic]. Interaction-based methods use context-response interactions to match potential responses [tao2021survey]. These methods followed a representation-matching-aggregation paradigm, formulating an interaction function to calculate the interaction between the two representation matrices of input utterances. The interaction function has two main types of definitions for the interaction function: similarity-based and attention-based methods [tao2021survey]. Similarity-based methods calculate the similarity of each word pair between the context message and the response candidate [wu2017sequential, zhang2018modeling, zhou2018multi]. Attention-based methods, however, use an attention mechanism to match the context message and the candidate’s response [chen2019sequential, humeau2019poly, yuan2019multi]. In recent years, pre-trained language models have been applied in retrieval-based dialogue systems due to their strong ability for language representation and understanding. These approaches employed an attention-based strategy to unify the representation, interaction, and aggregation operations by feeding the concatenation of context utterances and the candidate responses into a pre-trained multi-layer self-attention network [whang2020effective, gu2020speaker, xu2021learning, han2021fine, tao2021pre, feng2022reciprocal, li2022unsupervised].

### v3

@@ 7.1.2 Characteristics of e-commerce question answering
E-commerce QA services focus on answering product-aware questions asked by e-commerce users. Early studies on e-commerce QA focus on providing answers automatically from reviews by heuristic methods [li2009answering, moghaddam2011aqa, yu2012answering]. With the development of both QA techniques and e-commerce services, e-commerce QA has received increasing attention in recent years [mcauley2016addressing, yu2017modelling, yu2018aware, fan2019reading, zhang2020answer, gao2019product, gaoshen2021tois, feng2021multi, deng2022toward].
Distinct characteristics of e-commerce QA, as opposed to open-domain QA, are: (i) Domain-specific aspects are the first e-commerce QA characteristic. E-commerce QA systems rely on exploiting domain-specific information from product descriptions. Different products make different product-related aspects relevant or popular [mcauley2016addressing]. These product-aware aspects can help distinguish products and answer questions. (ii) There is a large number of consumer reviews, which can be used as a data source to help people form opinions and decisions [liu2016retrieving, mcauley2016addressing]. With the growth of those opinionated reviews, e-commerce users rely on advice from reviews before making purchase decisions. Reviews have been used as supporting data and candidate answers to supervise QA prediction models [yu2018aware]. (iii) There is a variety of answer sources. Most e-commerce QA services focus on extracting answers from reviews [mcauley2016addressing], and many e-commerce sites provide question answer pairs as knowledge bases for QA.
Text generation approaches have been studied to generate answers to given questions and reviews. Question reranking and answer reranking also have been studied [yu2017modelling, zhang2020answer]. E-commerce QA research can be divided into two directions: extractive product-aware QA and generative product-aware QA. The former focuses on extracting sentences or passages from reviews to answer questions, whereas the latter applies textual generation approaches to generate answers. We detail each type of e-commerce QA study in Section 7.1.3 and 7.1.4, respectively.
@@ 7.1.3 Extractive product-aware QA
Most e-commerce QA systems extract relevant sentences or fragments from the input text to answer the question given by the consumer. Early studies automatically extracted answers from reviews by heuristic unsupervised methods [li2009answering, moghaddam2011aqa]. Follow-up work mainly focuses on the matching between questions and reviews or candidate answers [mcauley2016addressing, yu2018aware]. yu2012answering proposed a framework for opinionated QA, which organizes reviews into a hierarchical structure and retrieves review sentences as the answer. The authors then use such a hierarchical structure to help retrieve questions and relevant review fragments. A joint optimization approach is proposed by simultaneously considering review salience, coherence, and diversity to rank fragments. liu2016retrieving find a concise set of questions addressed by a given review and cover its main points to help the user quickly comprehend the reviews. The authors propose a two-stage framework, where a probabilistic retrieval model is used to retrieve candidate questions and a matching procedure between answers and questions is used to bridge the vocabulary gap between reviews and questions.
@@
Some products, such as clothes and paintings, may not have proper names. Different strategies have been considered to replace the external knowledge of e-commerce to address this problem. mcauley2016addressing propose an answer prediction model by incorporating an aspect analytic model to learn latent aspect-specific review representation for predicting the answer. wan2016modeling address ambiguity, subjectivity, and diversity problems in consumer reviews. By using multiple answers in a supervised framework, the authors provide more accurate answers to objective and subjective questions. The authors also release a large-scale e-commerce QA dataset consisting of 135 thousand products from Amazon, 808 thousand questions, 3 million answers, and 11 million reviews. carmel2018product focus on subjective questions from Amazon customers, which can relate to various intent types such as product usage, recommendations, and opinions. The authors apply automatic QA methods, enhanced with community QA approaches to retrieve the most relevant answer found in reviews and QAs to address this problem. yu2018aware propose an answer prediction model by incorporating an aspect analytic model to learn latent aspect-specific review representation for predicting the answer. The authors establish the advantage of generating aspect-specific representations for new questions, which they use to develop a predictive answer model to capture intricate relationships among question texts and review texts. The proposed model uses reviews as a knowledge source to predict the answer by classifying answers into two types, binary (i.e. “yes” or “no”) and open-ended responses. As the amount of labeled data is limited in customer reviews, das2019learning propose an adversarial review-based approach to answer subjective and specific product-aware questions in a weakly supervised setting. Reading comprehension has been found to be useful to help extract relevant answers from e-commerce reviews [fan2019reading, xu2019review, zhang2020answering, chen2019answer]. Using the raw text of product-aware questions and customer reviews, fan2019reading introduce an end-to-end neural network model to synthesize multiple review representations. chen2019answer design a multi-task attentive model, namely QAR-Net, to identify plausible answers from product reviews for user questions. QAR-Net can use generated question answer pairs to help question-review matching.
@@
Pre-trained language models help understand the content of questions and reviews. xu2019review apply a BERT-based fine-tuning approach to extract answers from reviews. mittal2021distantly use a pre-trained language model to learn a relevance function by jointly learning unified syntactic and semantic representations of questions and reviews. A QA dataset for review comprehension with subjectivity labels for questions and answers has also been exploited [bjerva2020subjqa]. Besides user reviews, another type of information, namely product details provided by the manufacturer, has been considered an auxiliary information source for addressing product-related questions [zhang2020answering]. In order to alleviate the unavailability of labeled data, DBLP:conf/emnlp/JainRA23 introduce a distant supervision based model to prepare training data without manual effort.
@@
Review-based QA approaches extract answers from customer reviews, which can partially address users’ questions. However, there are many products with few or no reviews available. By collecting question answer pairs from real users, many e-commerce platforms develop retrieval-based QA systems for automatically answering frequently asked questions (FAQs) in the e-commerce industry [yu2017modelling, song2020tcnn, song2021online, zhang2020answer]. In Fig. 7.1, we see the retrieval-based QA framework applied by Alibaba [yu2017modelling]. Given a collection of question answer pairs (i.e., the knowledge base in Fig. 7.1), a key component is the question rerank module, which reranks candidate questions in a question answering knowledge base to find the best match given a question from a user. Based on such a framework, yu2017modelling formulate e-commerce QA as a paraphrase identification problem, where the target is to identify semantic relations of the given sentence pairs. The authors describe a transfer learning QA strategy to adapt the shared knowledge learned from a resource-rich source domain to a resource-poor target domain. Amazon has presented a large review-based QA dataset, namely AmazonQA, based on their real-world community QA platform [gupta2019amazonqa]. AmaonQA uses consumer reviews as the data resource and extracts snippets to answer questions. song2020tcnn improve the matching performance in retrieval-based e-commerce QA by introducing a multi-layer triple convolutional neural network model. Also, a sub-graph searching mechanism is shown to improve the efficiency of retrieval-based e-commerce QA [song2021online]. zhang2020answer focus on answer selection in retrieval-based e-commerce QA. Using graph neural networks, the authors jointly model multiple semantic relations, including semantic relevance between the question and answers, textual similarity among answers, and textual entailment between answers and reviews. rozen2021answering detail an answer prediction approach that uses similar questions about other products. The authors calculate contextual product similarity to determine whether two products are similar in the context of a specific question. Two large-scale datasets, including a question-to-question similarity dataset from Amazon and a corpus of question answer pairs from Amazon, have been released with the publication.
@@ 7.1.4 Generative product-aware QA
Many e-commerce portals have provided question answering services that assist users in posing product-aware questions to other consumers who have purchased the same product before. Users must read the product’s reviews to find the answer themselves. Given product attributes and reviews, following a cascading procedure, an answer is manually generated: (i) a user skims reviews and finds relevant sentences; (ii) they extract functional semantic units; and (iii) and the user jointly combines these semantic units with attributes and writes an appropriate answer. With a rapidly increasing number of reviews this process needs support [gao2019product]. Several strategies have been proposed to automatically generate answers using the product’s reviews to alleviate the burdens of customers [gao2019product, chen2019driven, deng2020opinion, lu2020chime, feng2021multi, deng2022toward]. The task on which these approaches focus is generative product-aware QA given reviews and product attributes.
@@
In first attempts, gao2019product, chen2019driven propose the task of product-aware answer generation, where a product-related question answering model is applied to incorporate customer reviews with product attributes. The authors formulate the research problem in generative e-commerce QA: for a product, there is a question $X^{q}=\{x^{q}_{1},x^{q}_{2},\dots,x^{q}_{T_{q}}\}$, $T_{r}$ reviews $X^{r}=\{x^{r}_{1},x^{r}_{2},\dots,x^{r}_{T_{r}}\}$ and $T_{a}$ key-value pairs of attributes $A=\{(a^{k}_{1},a^{v}_{1}),(a^{k}_{2},a^{v}_{2}),\dots,(a^{k}_{T_{a}},a^{v}_{T_{a}})\}$, where $a^{k}_{i}$ is the name of $i$-th attribute and $a^{v}_{i}$ is the attribute content. Each attribute, including key $a^{k}_{i}$ and value $a^{v}_{i}$, is represented as a single word in the generation task. Given a question $X^{q}$, an answer generator reads the reviews $X^{r}$ and attributes $A$, then generates an answer $\hat{Y}=\{\hat{y}_{1},\hat{y}_{2},\dots,\hat{y}_{T_{y}}\}$. The goal is to generate an answer $\hat{Y}$ that is grammatically correct and consistent with product attributes and opinions in the reviews.
@@
Fig. 7.2 provides an overview of the product-aware answer generator, the PAAG model, proposed by gao2019product. PAAG has four parts: (i) a review reader reads the review to extract relevant semantic parts; (ii) an attribute encoder encodes the attribute key-value pairs using a key-value memory network; (iii) a facts decoder generates the final answer according to the facts learned by the two modules introduced before; and (iv) a consistency discriminator distinguishes whether the generated answer matches the extracted facts, and we also use the result of the discriminator as another training signal. A generative e-commerce QA dataset extracted from JD.com is released with the publication. Similarly, chen2019driven formulate a noise-tolerant solution based on convolutional neural networks to generate natural answers.
@@
It is necessary to consider the text information from different reviews and attributes to answer specific questions in the wild. In Fig. 7.3, feng2021multi provide examples to demonstrate the multi-type text relation for product-aware question answering. As an example, Q1 asks “Does the design of this top look baggy?” R1 and R2 do not answer this question directly. But they provide a common entity “bat-like sleeve.” If we transfer the information provided by R1 and R2 to answer Q1 indirectly, it is easy to generate the answer that “The design of this tops looks baggy.” By integrating, understanding, and reasoning over the information of reviews and product attributes we may generate more accurate and pleasing answers to complex questions.
@@
A major limitation of most generative QA approaches is that they analyze each review and the corresponding attribute of the product individually, i.e., they neglect the relationship between different reviews/attributes of the product. feng2021multi propose a review-attribute heterogeneous graph neural network, RAHGNN, for product-aware answer generation to sufficiently understand and reason about the related information and its inner logic in multiple types of text. Most generative product-aware QA methods neglect personalization as it is insufficient to provide the same “completely summarized” answer to all customers. As an exception, deng2022toward describe a personalized answer generation method, PAGE, to model multi-perspective user preferences in personalized product question answering.
@@ 7.2 Dialogue systems in e-commerce
Dialogue systems have increasingly attracted attention in e-commerce. This section introduces studies on dialogue systems that can be applied to e-commerce platforms. Following previous work investigating this problem [Chen2017A, ziwwsdm1], we divide this section into three parts. We introduce recent studies on dialogue systems in Section 7.2.1, then detail task-oriented dialogue systems in e-commerce in Section 7.2.2, and discuss knowledge-grounded conversational agents in Section 7.2.3.
@@ 7.2.1 Introduction to dialogue systems
Dialogue systems are being considered in numerous applications, from e-commerce technical support to personal assistant tools [song2017summarizing, Chen2017A, ziwwsdm1, sun2016conversational, zhang2018towards, lei2018sequicity, liu2018knowledge, meng2020dukenet, sun2021conversations, shen2021vida, zhao2021jddc, liu2021conversational, ren2022variational, li2022knowledge, yu2022xdai]. The goal of creating an automatic human-computer conversational system as an assistant or chat companion is no longer an illusion now that two important factors have been seen progress. First, many conversation logs are now accessible, making it possible for machines to learn how to respond to input utterances. Second, deep generative neural network models, such as sequence-to-sequence and generative adversarial networks, are now able to capture complex patterns in large volumes of data [Chen2017A]. Based on these two factors, studies on dialogue systems focuses on methods to provide a natural and coherent response given an utterance from a user [Young2013POMDP, Ritter2011Data, Banchs2013IRIS, Ameixa2014Luke].
Dialogue systems can be divided into chitchat systems, task-oriented dialogue systems, and knowledge-grounded conversations [Chen2017A]. Chitchat agents are applied widely in open-domain dialogue systems, where dialogue systems interact with humans to provide reasonable and natural responses for open-domain dialogues [ziwwsdm1, yan2017building]. Chitchat messages usually represent user experiences and preferences, playing an essential role in many real-world applications. yan2017building reveal that most utterances in the online shopping scenario are chitchat messages.
Task-oriented dialogue systems aim to complete a specific task, e.g., restaurant reservation, along with a response generation process. Fig. 7.4 shows the four individual modules on which traditional task-oriented dialogue systems are based: natural language understanding, dialogue state tracking, policy learning, and natural language generation [wen2017latent, mrkvsic2015multi]. Given an utterance from a user, the system generates a proper response to address the user’s intention. In recent years, end-to-end task-oriented dialogue generation methods have been proposed to address the overall purpose more efficiently [wen2017network, lei2018sequicity, wu2019transferable].
@@
Knowledge-grounded conversations focus on generating a response with the correct knowledge to address the user’s utterance [meng2020dukenet]. Work on knowledge-grounded conversations can be categorized into two groups. Methods in the first group use structured knowledge (given knowledge graphs)  [zhou2018commonsense, liu2018knowledge, tuan2019dykgchat, wu2019proactive, moon2019opendialkg, wu2020diverse, zhou-etal-2020-kdconv, wang2020improving, wu2020improving, xu2020knowledge, xu2020conversational, jung2020attnio, xu-etal-2021-discovering]. Those in the second group use unstructured knowledge, such as document-based unstructured knowledge (given a whole document, e.g., a Wikipedia article) [chuanmeng2020refnet, ma2020survey, ma2020compare, tian-etal-2020-response, ren2019thinking, gopalakrishnan2019topical, zekangli2018incremental, qin-etal-2019-conversing, moghe2018towards, zhou2018dataset] or piece-based unstructured knowledge (given some separate pieces of knowledge, e.g., Foursquare tips) [ghazvininejad2018knowledge, dinan2018wizard, meng2020dukenet, kim2020sequential, Lian2019Learning, zheng2019enhancing, zheng2020difference, chen2020bridging, zheng2020approximation, zhao2020knowledge, yu2022xdai].
@@
With respect to generating responses, dialogue systems can be divided into retrieval-based and generation-based dialogue systems. The former retrieves several response candidates from a prebuilt index and then selects an appropriate one as a response. In contrast, the latter directly synthesizes a reply via natural language generation techniques [hvred, tao2021survey].
@@
Retrieval-based dialogue systems retrieve several response candidates from a prebuilt index and then select an appropriate one as a response. Social networks have accumulated a significant amount of conversational data among humans on the web, motivating researchers to investigate data-driven approaches to re-use human conversations and select a response for new input from candidates [tao2021survey, xu2021response]. Retrieval-based dialogue generation methods outperform their generation-based counterparts in response fluency and informativeness. They power a series of real-world applications, e.g., XiaoIce from Microsoft [zhou2020design]. Learning to rank and matching approaches have been widely applied in the retrieval process [yan2016learning, wu2017sequential, zhou2018multi, yang2018response, yuan2019multi, su2020dialogue, tao2021survey, lin2022task]. A core task in retrieval-based dialogue systems is response selection. Studies into retrieval-based response selection can be divided into three types: representation-based, interaction-based, and pre-trained language model-based methods. Representation-based methods are composed of a representation-matching paradigm and consist of a representation layer and a matching layer [yan2016learning, wu2018response, wang2017deep, zhou2018dataset, zhou2018multi, yan2018response, xu2021topic]. Interaction-based methods use context-response interactions to match potential responses [tao2021survey]. These methods follow a representation-matching-aggregation paradigm, formulating an interaction function to calculate the interaction between the two representation matrices of input utterances. The interaction function has two main types of definition: similarity-based and attention-based methods [tao2021survey]. Similarity-based methods calculate the similarity of each word pair between the context message and the response candidate [wu2017sequential, zhang2018modeling, zhou2018multi]. Attention-based methods, however, use an attention mechanism to match the context message and the candidate’s response [chen2019sequential, humeau2019poly, yuan2019multi]. In recent years, pre-trained language models have been applied in retrieval-based dialogue systems due to their strong ability for language representation and understanding. These approaches employ an attention-based strategy to unify the representation, interaction, and aggregation operations by feeding the concatenation of context utterances and the candidate responses into a pre-trained multi-layer self-attention network [whang2020effective, gu2020speaker, xu2021learning, han2021fine, tao2021pre, feng2022reciprocal, li2022unsupervised].